# Hypergraphs as Weighted Directed Self-Looped Graphs: Spectral Properties, Clustering, Cheeger Inequality

### v1

@@ 2 Preliminaries
$R$ ||×$|\mathcal{E}|\times|\mathcal{V}|$ vertex-weight matrix
@@
A hypergraph consists of vertices and hyperedges. A hyperedge $e$ is a connection between two or more vertices. We use the notation $v\in e$ if the hyperedge $e$ connects vertex $v$. This is also called “$e$ is incident to $v$”. We first provide the formal definition of an EDVW hypergraph. Definition 1 and 2 provide necessary notations to define the hypergraph random walk in Definition 3. The transition matrix $P$ of EDVW hypergraphs is consistent with that of graphs.
@@
[7] (EDVW hypergraph). A hypergraph $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ with edge-dependent vertex weight is defined as a set of vertices $\mathcal{V}$, a set $\mathcal{E}\subseteq 2^{\mathcal{V}}$ of hyperedges, a weight mapping $\omega(e):\mathcal{E}\rightarrow\mathbb{R}_{+}$ on every hyperedge $e\in\mathcal{E}$, and weight mappings $\gamma_{e}(v):\mathcal{V}\rightarrow\mathbb{R}_{\geq 0}$ corresponding to $e$ on every vertex $v$. For $e_{1}\neq e_{2}$, $\gamma_{e_{1}}(v)$ and $\gamma_{e_{2}}(v)$ may be different. Without loss of generality, we index the vertices by $1,2,...,|\mathcal{V}|$, and let $\mathcal{V}=\{1,2,...,|\mathcal{V}|\}$.
@@
[7] (Vertex-weight matrix, hyperedge-weight matrix, vertex-degree matrix and hyperedge-degree matrix of an EDVW hypergraph). $E(v)=\{e\in\mathcal{E}\,s.t.\,v\in e\}$ is the set of hyperedges incident to vertex $v$. $d(v)=\sum_{e\in E(v)}w(e)$ denotes the degree of vertex $v$. $\delta(e)=\sum_{v\in e}\gamma_{e}(v)$ denotes the degree of hyperedge $e$. The vertex-weight matrix $R$ is an ||×$|\mathcal{E}|\times|\mathcal{V}|$ matrix with entries $R(e,v)=\gamma_{e}(v)$. The hyperedge-weight matrix $W$ is a $|\mathcal{V}|\times|\mathcal{E}|$ matrix with entries $W(v,e)=\omega(e)$ if $v\in e$, and $W(v,e)=0$ otherwise. The vertex-degree matrix $D_{\mathcal{V}}$ is a $|\mathcal{V}|\times|\mathcal{V}|$ diagonal matrix with entries $D_{\mathcal{V}}=d(v)$. The hyperedge-degree matrix $D_{\mathcal{E}}$ is a $|\mathcal{E}|\times|\mathcal{E}|$ diagonal matrix with entries $D_{\mathcal{E}}(e,e)=\delta(e)$.
@@
Table 2 contains important notation and hyperparameters for quick reference. The random walk on hypergraph with edge-dependent vertex weights was proposed in [7]. Intuitively, at time $t$, a random walker at vertex $u$ will first pick an edge $e$ incident to $u$ with the probability $\frac{\omega(e)}{d(u)}$, then pick a vertex $v$ from the picked edge $e$ with the probability $\frac{\gamma_{e}(v)}{\delta(e)}$, and finally move to vertex $v$ at time $t+1$. Define the transition matrix $P$ to be a $|\mathcal{V}|\times|\mathcal{V}|$ matrix with entries $P_{u,v}$ to be the transition probability from $u$ to $v$.
@@
[7] (Hypergraph random walk). A random walk on a hypergraph with edge-dependent vertex weights $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ is a Markov Chain on $\mathcal{V}$ with transition probabilities
@@
$\phi P=\phi;\leavevmode\nobreak\ \phi(u)>0\,\,\forall u\in\mathcal{V};\leavevmode\nobreak\ \sum_{u\in{\mathcal{V}}}\phi(u)=1$ (5)
From $\phi$, we further define the stationary distribution matrix to be a $|\mathcal{V}|\times|\mathcal{V}|$ diagonal matrix with entries $\Pi_{i,i}=\phi(i)$.
@@
[7]. The stationary distribution $\phi$ of hypergraph random walk exists.
Theorem 5 has been proved in [7]. In this paper, we give a simplified proof in Appendix A.4.
@@
[7] (random-walk-based hypergraph Laplacian). Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be a hypergraph with edge-dependent vertex weight. Let $P$ be the transition matrix and $\Pi$ be the corresponding stationary distribution matrix. Then, the random-walk-based hypergraph Laplacian $L$ is
@@ 3 Spectral Properties Inspire Global Partitioning
In this section, we first extend the spectral theory for EDVW hypergraphs, then introduce our HyperClus-G algorithm. Specifically, we show that using our definitions, there are algebraic connections between hypergraph NCut, Rayleigh Quotient and Laplacian (Theorem 1), which are consistent to pair-wise graphs.
@@
$\begin{split}\phi&(S)=(\phi P)(S)\\ &=\phi(S)-\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}+\sum_{u\in\bar{\mathcal{S}},v\in\mathcal{S}}\phi(u)P_{u,v}\end{split}$ (8)
@@
For unweighted and undirected graphs, the volume of the boundary/cut of a partition is defined as $|\partial\mathcal{S}|=|\{\{x,y\}\in E|x\in\mathcal{S},y\in\bar{\mathcal{S}}\}|$. The intuition behind this definition is the symmetric property $|\partial\mathcal{S}|=|\partial\bar{\mathcal{S}}|$. Theorem 6 also describes such a property, and we find it intuitively suitable to be extended to the following definition.
@@
According to the form of generalized Rayleigh Quotient $R_{L}(x)=\frac{x^{T}Lx}{x^{T}Dx}$ [19], we extend this generalized Rayleigh Quotient to hypergraphs as follows.
@@
Proof.
Notice that when $x^{T}Ax$ is a scalar,
$x^{T}Ax=(x^{T}Ax)^{T}=((x^{T}A)x)^{T}=x^{T}(x^{T}A)^{T}=x^{T}A^{T}x$ (15)
we have
$x^{T}\Pi Px=x^{T}P^{T}\Pi^{T}x=x^{T}P^{T}\Pi x$ (16)
[…]
The 2-way Normalized Cut, a well-known measurement of the cluster quality, is defined as $\frac{|\partial\mathcal{S}|}{vol(\mathcal{S})}+\frac{|\partial\mathcal{\bar{S}}|}{vol(\bar{\mathcal{S}})}$. We derive the NCut for EDVW hypergraphs using Definition 7.
@@
Following this definition, the problem of global partitioning on EDVW hypergraphs aims to find a vertex set $\mathcal{S}$ that minimizes the NCut value. We now derive the association between our hypergraph Ncut, Rayleigh Quotient, and Laplacian. Such association directly provides the inspiration for a spectral 2-way clustering approach (Algorithm 1).
@@
For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, we define a $|\mathcal{V}|$-dimensional vector $x$ such that
$\begin{split}x(u)&=\sqrt{\frac{vol(\bar{\mathcal{S}})}{vol(\mathcal{S})}},\forall\,\,u\in\mathcal{S},\\ x(\bar{u})&=-\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}},\forall\,\,\bar{u}\in\bar{\mathcal{S}}.\\ \end{split}$ (19)
@@
Proof.
$x(u)=\begin{cases}\sqrt{\frac{vol(\mathcal{S})}{vol(\mathcal{S})}}&\text{if }u\in\mathcal{S}\\ -\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}}&\text{if }u\in\bar{\mathcal{S}}\end{cases}$ (21)
From the definition of $NCut$ in Eq 10, we can compute $R(x)$ as follows:
@@
The above is an NP-complete problem [39]. If we relax the restriction of $x$ that each entry has to be either $\sqrt{\frac{vol(\bar{\mathcal{S}})}{vol(\mathcal{S})}}$ or $-\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}}$ for some $\mathcal{S}$, then the problem becomes
$\min_{x}R(x)\Leftrightarrow\min_{x}\frac{x^{T}Lx}{x^{T}\Pi x},\leavevmode\nobreak\ \text{for }x\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}.$ (24)
Using the property of Rayleigh Quotient, the minimum can be achieved by choosing $x$ to be the eigenvector associated with the second smallest eigenvalue of the generalized eigenvalue system $Lx=\lambda\Pi x$. Replacing $z=\Pi^{\frac{1}{2}}x$, we have
@@
$\Pi^{-\frac{1}{2}}L\Pi^{-\frac{1}{2}}z=\lambda_{1}z$ (26)
$\begin{cases}u\in S&\text{ if }z(u)\geq 0\\ u\in\bar{S}&\text{ if }z(u)<0\par\end{cases}$ (27)
Moreover, we define a new variant of hypergraph Laplacian.
Definition 11.
$L_{sym}$ is also real and symmetric, and hence hermitian. The eigenvectors of $L_{sym}$ provide an approximate solution of 2-way hypergraph clustering. The formulation of this Laplacian is consistent with that for digraphs [8].
4 Hypergraph Cheeger Inequality and Optimality of HyperClus-G
In this section, we show the hypergraph Cheeger Inequality (Theorem 3), which is consistent of the Cheeger Inequality of pair-wise graphs. The Cheeger Inequality requires the below definition of conductance. From the cheeger inequality, we prove the approximate linear optimality of HyperClus-G in terms of both NCut and Conductance.
Definition 12.
@@ 4 Hypergraph Cheeger Inequality and Optimality of HyperClus-G
For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, our hypergraph conductance $\Phi(\mathcal{S})\in[0,1]$, which is consistent with graph conductance
Proof.
Recall from the definition 12, 7 and 8,
$\Phi(\mathcal{S})=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}=\frac{\sum_{u\in\mathcal{S},v\in\mathcal{\bar{\mathcal{S}}}}\phi(u)P_{u,v}}{\min(\sum_{v\in\mathcal{S}}\phi(u),\sum_{u\in\bar{\mathcal{S}}}\phi(u))}$ (30)
From the definition of $P_{u,v}$ and $\phi(u)$, $\Phi(\mathcal{S})$ is non-negative.
[…]
Thus, $\min(\sum_{v\in\mathcal{S}}\phi(u),\sum_{u\in\bar{\mathcal{S}}}\phi(u))\geq\sum_{u\in\mathcal{S},v\in\mathcal{\bar{\mathcal{S}}}}\phi(u)P_{u,v}$ and $\Phi(\mathcal{S})\in[0,1]$. ∎
4.1 Hypergraph Cheeger Inequality
Cheeger Inequality can be proved for our newly defined Symmetric normalized random-walk-based hypergraph Laplacian (Definition 11). The proof is similar to that for digraphs [8] and shows the rationality of our definitions.
Theorem 10.
(Hypergraph Cheeger Inequality) Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be any hypergraph in the EDVW formatting with positive edge weights $\omega(\cdot)>0$ and non-negative edge-dependent vertex weights $\gamma_{e}(\cdot)$ for any $e\in\mathcal{E}$. Define $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$, where $\Phi(\mathcal{S})$ is defined in Definition 12. Then the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $L_{sym}$ satisfies
@@ 4.1 Hypergraph Cheeger Inequality
Proof.
From Theorem 14,
Thus the $\leq$” side has been proved. For the $\geq$” side, assume $w$ is the eigenvector of the normalized Laplacian associated with $\lambda$, let vector $f=\Pi^{-\frac{1}{2}}w$. Then,
Therefore, $\forall u$
$\begin{split}\lambda&=\frac{\sum_{u\in\mathcal{V}_{+}}f(u)\sum_{v}(f(u)-f(v))(\phi(v)P_{v,u}+\phi(u)P_{u,v})}{\sum_{u\in\mathcal{V}_{+}}f(u)2f(u)\phi(u)}\\ &=\frac{\sum_{u\in\mathcal{V}}g(u)\sum_{v}(f(u)-f(v))(\phi(v)P_{v,u}+\phi(u)P_{u,v})}{\sum_{u\in\mathcal{V}}2g(u)^{2}\phi(u)}\\ &\geq\frac{\sum_{u}g(u)\sum_{v}(g(u)-g(v))(\phi(v)P_{v,u}+\phi(u)P_{u,v})}{\sum_{u}2g(u)^{2}\phi(u)}\\ \end{split}$ (37)
One can validate that
Continue with Equation 37 with Theorem 3 in [8],
$\begin{split}\lambda&\geq\frac{\sum_{u}\sum_{v}(g(u)-g(v))^{2}\phi(v)P_{v,u}}{\sum_{u}2g(u)^{2}\phi(u)}\\ &=\frac{(\sum_{u}\sum_{v}(g(u)-g(v))^{2}\phi(v)P_{v,u})^{2}}{\sum_{u}2g(u)^{2}\phi(u)(\sum_{u}\sum_{v}(g(u)-g(v))^{2}\phi(v)P_{v,u})}\\ &\geq\frac{(\sum_{u}\sum_{v}|g(u)^{2}-g(v)^{2}|\phi(v)P_{v,u})^{2}}{8(\sum_{u}g(u)^{2}\phi(u))^{2}}\\ &=\frac{(\sum_{k}\sum_{l>k}(g(v_{k})^{2}-g(v_{l})^{2})(\phi(v_{l})P(v_{l},v_{k})+\phi(v_{k})P(v_{k},v_{l})))^{2}}{8(\sum_{u}g(u)^{2}\phi(u))^{2}}\\ &=\frac{(\sum_{k}(g(v_{k})^{2}-g(v_{k+1})^{2})\sum_{i\leq k<j}(\phi(v_{i})P(v_{i},v_{j})+\phi(v_{j})P(v_{j},v_{i})))^{2}}{8(\sum_{u}g(u)^{2}\phi(u))^{2}}\\ \end{split}$ (40)
Using Equation 38, continue with Equation 40,

### v2

@@ 2 Preliminaries
$R$ $|\mathcal{E}|\times|\mathcal{V}|$ vertex-weight matrix
@@
Table 3 contains important notation and hyperparameters for quick reference. A hypergraph consists of vertices and hyperedges. A hyperedge $e$ is a connection between two or more vertices. We use the notation $v\in e$ if the hyperedge $e$ connects vertex $v$. This is also called “$e$ is incident to vv". Definition 4 and 5 provide necessary notations to define the hypergraph random walk in Definition 6. The transition matrix $P$ of EDVW hypergraphs is consistent with that of graphs.
@@
(Chitra & Raphael, 2019) (EDVW hypergraph). A hypergraph $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ with edge-dependent vertex weight is defined as a set of vertices $\mathcal{V}$, a set $\mathcal{E}\subseteq 2^{\mathcal{V}}$ of hyperedges, a weight mapping $\omega(e):\mathcal{E}\rightarrow\mathbb{R}_{+}$ on every hyperedge $e\in\mathcal{E}$, and weight mappings $\gamma_{e}(v):\mathcal{V}\rightarrow\mathbb{R}_{\geq 0}$ corresponding to $e$ on every vertex $v$. For $e_{1}\neq e_{2}$, $\gamma_{e_{1}}(v)$ and $\gamma_{e_{2}}(v)$ may be different. Without loss of generality, we index the vertices by $1,2,...,|\mathcal{V}|$, and let $\mathcal{V}=\{1,2,...,|\mathcal{V}|\}$.
@@
(Chitra & Raphael, 2019) (Vertex-weight matrix, hyperedge-weight matrix, vertex-degree matrix, and hyperedge-degree matrix of an EDVW hypergraph). $E(v)=\{e\in\mathcal{E}\,s.t.\,v\in e\}$ is the set of hyperedges incident to vertex $v$. $d(v)=\sum_{e\in E(v)}w(e)$ denotes the degree of vertex $v$. $\delta(e)=\sum_{v\in e}\gamma_{e}(v)$ denotes the degree of hyperedge $e$. The vertex-weight matrix $R$ is an $|\mathcal{E}|\times|\mathcal{V}|$ matrix with entries $R(e,v)=\gamma_{e}(v)$. The hyperedge-weight matrix $W$ is a $|\mathcal{V}|\times|\mathcal{E}|$ matrix with entries $W(v,e)=\omega(e)$ if $v\in e$, and $W(v,e)=0$ otherwise. The vertex-degree matrix $D_{\mathcal{V}}$ is a $|\mathcal{V}|\times|\mathcal{V}|$ diagonal matrix with entries $D_{\mathcal{V}}=d(v)$. The hyperedge-degree matrix $D_{\mathcal{E}}$ is a $|\mathcal{E}|\times|\mathcal{E}|$ diagonal matrix with entries $D_{\mathcal{E}}(e,e)=\delta(e)$.
@@
The random walk on EDVW hypergraph was proposed in (Chitra & Raphael, 2019). Intuitively, at time $t$, a random walker at vertex $u$ will first pick an edge $e$ incident to $u$ with the probability $\frac{\omega(e)}{d(u)}$, then pick a vertex $v$ from the picked edge $e$ with the probability $\frac{\gamma_{e}(v)}{\delta(e)}$, and finally move to vertex $v$ at time $t+1$. We can define the transition matrix $P$ to be a $|\mathcal{V}|\times|\mathcal{V}|$ matrix with entries $P_{u,v}$ to be the transition probability from $u$ to $v$ as follows.
@@
(Chitra & Raphael, 2019) (Hypergraph random walk). A random walk on a hypergraph with edge-dependent vertex weights $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ is a Markov Chain on $\mathcal{V}$ with transition probabilities
@@
Random walks and stationary distribution serve as a basis when conducting spectral analysis. Under assumption 1, a unique stationary distribution exists (referring to Theorem 8).
@@
$\phi P=\phi;~\phi(u)>0\,\,\forall u\in\mathcal{V};~\sum_{u\in{\mathcal{V}}}\phi(u)=1$ (5)
From $\phi$, define the stationary distribution matrix to be a $|\mathcal{V}|\times|\mathcal{V}|$ diagonal matrix with entries $\Pi_{i,i}=\phi(i)$.
@@
(Chitra & Raphael, 2019). The stationary distribution $\phi$ of hypergraph random walk exists.
Theorem 8 has been proved in (Chitra & Raphael, 2019). In this paper, we give a simplified proof in Appendix A.4.
@@
(Chitra & Raphael, 2019) (random-walk-based hypergraph Laplacian). Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be a hypergraph with edge-dependent vertex weight. Let $P$ be the transition matrix and $\Pi$ be the corresponding stationary distribution matrix. Then, the random-walk-based hypergraph Laplacian $L$ is
@@ 3 Spectral Properties Inspire Global Partitioning
In this section, we first extend the spectral theory for EDVW hypergraphs, then we introduce our HyperClus-G algorithm. Specifically, we show that using our definitions, there are algebraic connections among hypergraph NCut, Rayleigh Quotient, and Laplacian (Theorem 1), which are consistent with pair-wise graphs.
@@
$\begin{split}\phi&(S)=(\phi P)(S)=\phi(S)-\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}+\sum_{u\in\bar{\mathcal{S}},v\in\mathcal{S}}\phi(u)P_{u,v}\end{split}$ (8)
@@
For unweighted and undirected graphs, the volume of the boundary/cut of a partition is defined as $|\partial\mathcal{S}|=|\{\{x,y\}\in E|x\in\mathcal{S},y\in\bar{\mathcal{S}}\}|$. The intuition behind is the symmetric property $|\partial\mathcal{S}|=|\partial\bar{\mathcal{S}}|$. Theorem 11 also describes such a property, and we find it intuitively suitable to be extended to the following definition.
@@
According to the form of generalized Rayleigh Quotient $R_{L}(x)=\frac{x^{T}Lx}{x^{T}Dx}$ (Golub & Loan, 1996), we extend this generalized Rayleigh Quotient to EDVW hypergraphs as follows.
@@
(Proof in Appendix A.6)
The 2-way Normalized Cut, a well-known measurement of the cluster quality, is defined as $\frac{|\partial\mathcal{S}|}{vol(\mathcal{S})}+\frac{|\partial\mathcal{\bar{S}}|}{vol(\bar{\mathcal{S}})}$. Analogically, we derive the NCut for EDVW hypergraphs using Definition 12.
@@
Following this definition, the problem of global partitioning on EDVW hypergraphs aims to find a vertex set $\mathcal{S}$ that minimizes the NCut value. We now derive the association between our hypergraph Ncut, Rayleigh Quotient, and Laplacian, which directly inspires a spectral 2-way clustering approach (Algorithm 1).
@@
(Restatement of Theorem 1 in Main Results) For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, we define a $|\mathcal{V}|$-dimensional vector $x$ such that
[…]
@@
(Proof in Appendix A.7)
@@
The above is an NP-complete problem (Shi & Malik, 2000). If we relax the restriction of $x$ that each entry has to be either $\sqrt{\frac{vol(\bar{\mathcal{S}})}{vol(\mathcal{S})}}$ or $\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}}$ for some $\mathcal{S}$, then the problem becomes
$\min_{x}R(x)\Leftrightarrow\min_{x}\frac{x^{T}Lx}{x^{T}\Pi x},~\text{for }x\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}.$ (19)
Using the property of Rayleigh Quotient, the minimum can be achieved by choosing $x$ to be the eigenvector associated with the second smallest eigenvalue of the generalized eigenvalue system $Lx=\lambda\Pi x$ Ghojogh et al. (2019). Replacing $z=\Pi^{\frac{1}{2}}x$, we have
@@
$\Pi^{-\frac{1}{2}}L\Pi^{-\frac{1}{2}}z=\lambda_{1}z$ (21)
[…]
Moreover, we define a new variant of hypergraph Laplacian.
Definition 17.
$L_{sym}$ is also real and symmetric, and hence hermitian. The eigenvectors of $L_{sym}$ provide an approximate solution of 2-way hypergraph clustering. The formulation of this Laplacian is consistent with that for digraphs (Chung, 2005).
4 Hypergraph Cheeger Inequality and Optimality of HyperClus-G
In this section, we show the hypergraph Cheeger Inequality (Theorem 3), which is consistent with the Cheeger Inequality of pair-wise graphs. The Cheeger Inequality requires the below definition of conductance. From the Cheeger Inequality, we prove the approximate linear optimality of HyperClus-G in terms of both NCut and Conductance.
Definition 18.
Theorem 19.
For any vertex set $\mathcal{S}\subseteq\mathcal{V}$, our hypergraph conductance $\Phi(\mathcal{S})\in[0,1]$, which is consistent with graph conductance (Proof in Appendix A.8)
@@ 4 Hypergraph Cheeger Inequality and Optimality of HyperClus-G
4.1 Hypergraph Cheeger Inequality
Cheeger Inequality can be proved for our newly defined Symmetric normalized random-walk-based hypergraph Laplacian (Definition 17). The proof is similar to that for digraphs (Chung, 2005) and shows the rationality of our definitions.
@@ 4.1 Hypergraph Cheeger Inequality
(Hypergraph Cheeger Inequality, Restatement of Theorem 3 in Main Results) Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\omega,\gamma)$ be an EDVW hypergraph. Define $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$, where $\Phi(\mathcal{S})$ is defined in Definition 18. Then the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $L_{sym}$ satisfies
@@
(Proof in Appendix A.9)